OpenStudy (anonymous):

Mona has a triangular flag. She wants to cut out a similar flag that has dimensions one-fourth of the original flag. If the area of the original flag is 320 square inches, what is the area of the new flag that Mona cuts out? Answer 16 in2 20 in2 32 in2 80 in2

OpenStudy (experimentx):

20

OpenStudy (anonymous):

i tried deviding 320 by a fourth and it didnt work well it didnt give me an answer that was one of my choices at least

OpenStudy (anonymous):

if the dimensions (1 dimensional lengths are 1/4 of the original) then the area (2 dimensions) is (1/4)^2 smaller (power 2 because area is 2 dimensions) (1/4)^2 = 1/16 1/16 of 320 is 20

OpenStudy (experimentx):

(1/4)^ = 1/16